what is -5.00A + 4.00B

1
Expert's answer
2020-10-07T10:11:43-0400

For this question angle and co-ordinate points must be given for A and B.

Let the co-ordinate of A is"(x_1, y_1)" and angle with positive x axis is"= \theta"

And co-ordinate for the point B is "(x_2,y_2)" and it is making an angle "\phi"

So, "A=x_1 \cos\theta \hat{i}+y_1\sin\theta\hat{j}"

Similarly, "B=x_2\cos\phi\hat{i}+y_2\sin\phi\hat{j}"

Now it is given "C=-5.00A+4.00B"

Now, substituting the values,

"\Rightarrow C=-5(x_1 \cos\theta \hat{i}+y_1\sin\theta\hat{j})+4.00(x_2\cos\phi\hat{i}+y_2\sin\phi\hat{j})"

"\Rightarrow C=(4x_2\cos\phi-5x_1 \cos\theta)\hat{i}+(4y_2\sin\phi-5y_1\sin\theta)\hat{j})"